OpenStudy (anonymous):

MEDAL AWARD Help !! Find the equation of a line in slope-intercept form that passes through the points (-2, 20) and (3, -10).

OpenStudy (whpalmer4):

The equation for the slope (\(m\)) of a line passing passing through points (\(x_1,y_1\)), (\(x_2,y_2\)) is \[m=\frac{y_2-y_1}{x_2-x_1}\] Once you've determined the slope, you can write the equation for the line using point-slope form: \[y-y_1 = m(x-x_1)\]and then rearrange it into slope-intercept form:\[y=mx+b\]

OpenStudy (anonymous):

is y2 3? and y1 -10, x2 20 and x1 -2 ??

OpenStudy (whpalmer4):

You can choose \((x_1,y_1) = (-2,20)\) or \((x_1, y_1) = (3,-10)\)

OpenStudy (whpalmer4):

and hopefully it is obvious that the other point will be \((x_2,y_2)\)

OpenStudy (anonymous):

so either one you choose u get the same answer ?

OpenStudy (whpalmer4):

exactly.

OpenStudy (anonymous):

are u supposed to get 22/-13 for the first part

OpenStudy (whpalmer4):

no.

OpenStudy (whpalmer4):

show me your work

OpenStudy (anonymous):

-10-20/3--2 I got -30/5

OpenStudy (anonymous):

???

OpenStudy (whpalmer4):

that's right, the slope is -6. If you plot the two points, one is slightly to the left of the y-axis, and high. the other point is slightly to the right of the y-axis, and much lower. For each point you move to the right on the x-axis, you drop 6 points on the y-axis. That's what a slope of -6 means.